# Maximal Extractable Value in Decentralized Finance: Taxonomy, Detection, and Mitigation

### v1

@@ Abstract
Decentralized Exchanges (DEXes) are one of the prominent applications of DeFi, enabling users to directly swap tokens using smart contracts [2]. Most DEXes operate on the Ethereum blockchain to facilitate trustless and automated transactions. However, in the Ethereum network, transactions are stored in a public mempool before they are included in a block. This creates profitable opportunities for network participants, referred to as searchers, who can submit new transactions by observing pending financial transactions to gain additional revenue. These searchers maximize their profits by manipulating gas prices to influence the order of their transactions within the block. This additional value extracted from the blockchain network is termed Maximal Extractable Value (MEV). MEV searchers can be the block producers (i.e., miners in Proof of Work (PoW) and validators in Proof of Stake (PoS)), other network participants, or a bot. However, block producers have a unique advantage, as they can include transactions in blocks without paying high gas prices [3]. Furthermore, the transition from PoW to PoS in Ethereum has significantly reduced block producer rewards, further luring block producers to engage in MEV activities [4].
@@
Several surveys have explored various aspects of MEV, including transaction types and mitigation strategies [7, 8, 6, 5, 9]. In particular, [7] classifies MEV transactions, while [8, 6] emphasize on mitigation techniques. In contrast, [5, 9] cover both transaction types and mitigation strategies. However, none have thoroughly examined all transaction types, detection approaches, and mitigation strategies. This comprehensive survey addresses this gap by presenting a novel taxonomy of MEV transaction types supported by real-world examples from the Ethereum network. It also explores MEV detection approaches and mitigation strategies and examines various simulation and extraction methods. The main contributions of this survey are as follows.
We introduce a novel and comprehensive taxonomy of MEV transactions, supported by real Ethereum transaction examples for each identified MEV type. The taxonomy clearly distinguishes between value-diverting MEV transactions, which can lead to financial loss and network instability, and value-creating MEV transactions, which can stabilize markets or enhance efficiency. This distinction provides clarity on the dual nature of MEV activities and lays the foundation for discussions on mitigation and detection strategies.
@@
We explore existing MEV simulation frameworks and extraction methods, offering a comprehensive overview of how these tools model and replicate real-world MEV scenarios.
We critically examine the major challenges in MEV mitigation and detection, such as centralization, latency, layer-2 MEV, and multi-address MEV. Furthermore, we discuss potential solutions paving the way for a more secure and efficient DeFi system.
@@ 2.1.1 EIP-1599
2.2 Decentralized Exchanges (DEXes)
In traditional centralized exchanges, a central exchange operator maintains a list of assets trade offers from different buyers and sellers [14]. A seller’s order specifies the minimum price at which they are willing to sell an asset, whereas a buyer’s order specifies the maximum price at which they are willing to buy an asset [11]. The operator triggers balance changes after matching the sell-buy orders. In contrast, Decentralized Exchanges (DEXes) eliminates the need for a centralized operator [2]. Instead, a smart contract is used to match the buyer and seller options. The list of buy and sell orders is maintained off-chain. Order matching is performed either by the trader (e.g., in Etherdelta) or by the smart contract (e.g., in IDex) [11].
2.2.1 Automated Market Makers (AMMs)
DEXes implemented within blockchain use a mechanism that enables asset trading without relying on traditional order books of buy and sell orders. This mechanism is known as Automated Market Makers (AMMs), and it utilizes mathematical models to determine asset exchange prices and execute trades. AMMs manage reserves of assets in what are called liquidity pools, where liquidity providers deposit pairs of assets. When a trade occurs, the AMM adjusts the price for subsequent trades based on the current ratio of assets in the liquidity pool. Larger trades can lead to slippage, which is a significant change in the asset’s exchange price.
AMMs utilize various mathematical models to determine how token prices are set within their liquidity pools. For a comprehensive overview of various AMM models and their mechanisms, we encourage readers to refer to the work of Xu et al. [2].Buying (selling) a token reduces (increases) its quantity in the liquidity pool, causing its exchange price to rise (fall).
@@ 2.3 Maximal Extractable Value (MEV)
Public transactions on Ethereum enable value extraction through the insertion, removal, or reordering of transactions [11]. The acquired value is known as the Maximal Extractable Value (MEV). According to Ethereum’s documentation444https://ethereum.org/en/developers/docs/mev/, accessed on 05 September 2024: Maximal extractable value (MEV) refers to the maximum value that can be extracted from block production in excess of the standard block reward and gas fees by including, excluding, and changing the order of transactions in a block.
@@
Network participants who seek profitable opportunities for value extraction are called searchers. This search for MEV can be conducted either by human users or by automated bots. Therefore, the terms searchers and MEV bots are often used interchangeably. Before the Merge555The update that transitioned Ethereum’s consensus protocol from Proof-of-Work (PoW) to Proof-of-Stake (PoS) to improve scalability, security, and energy efficiency, Ethereum was based on the Proof-of-Work (PoW) consensus protocol, where block producers were known as miners. The value extracted from block production was thus initially referred to as Miner Extractable Value. After the Merge in September 2022, Ethereum transitioned to the Proof-of-Stake (PoS) consensus protocol, with validators taking the place of miners. Additionally, a searcher does not necessarily have to be the block producer; they can also be non-miner users or bots. Consequently, the term ”Miner” in MEV was replaced with ”Maximal” to reflect this change. MEV is sometimes also referred to as ”Blockchain Extractable Value” [15]. Furthermore, the Flashbots project refers to the actual extracted values due to MEV opportunities as ”Realized Extractable Value”666https://writings.flashbots.net/quantifying-rev, accessed on 05 September 2024. Judmayer et al. [16] describe the potential or future values that could be extracted if a sequence of transactions eventually becomes a part of the blockchain (with some probability) as ”Expected Extracted Value”.
@@
Searchers in the Ethereum network often capitalize on price slippage in DEXes to acquire MEV. However, not all MEV transactions can be considered attacks. Some searchers may monitor the mempool or DEXes and execute MEV transactions that can actually improve the system’s efficiency. For example, a searcher may identify a price difference for the same asset across different DEXes (this arbitrage opportunity arises due to buying or selling activities). The searcher performs an MEV transaction to profit from this price difference, which helps stabilize asset prices across the DEXes. Such transactions are known as ”value-creating” MEV transactions, as they generate new value [6]. These transactions are generally not considered attacks.
@@
On the other hand, some searchers may monitor the mempool for transactions aimed at acquiring MEV and replace them with their own transactions, or they might create a transaction that generates an MEV opportunity and later execute another transaction to capture that MEV. For example, a searcher might detect a transaction aimed at exploiting an arbitrage opportunity, copy it, and then resubmit it with a higher gas fee to capture the MEV before the original transaction is executed. Such transactions are referred to as ”value-diverting” MEV transactions because they divert value from other users to the searcher [6]. Such transactions can be considered attacks, as they negatively impact other users. Value-diverting transactions have a detrimental impact on the DeFi system, especially when multiple searchers engage in open-bid first-price777In a first-price auction, the winning searcher pays their bid amount. on-chain auctions to capture MEV. These competitive auctions, known as Priority Gas Auctions (PGAs) [11], force searchers to pay significantly higher gas fees to prioritize their transactions. This bidding war increases the average gas price across the network, discouraging honest users from utilizing the platform. Additionally, searchers often submit multiple redundant transactions with progressively higher gas fees to maintain their position in the auction, leading to network congestion and increased latency [6]. Daian et al. [11] illustrates an example of a PGA between two searchers, where the winning searcher issued 42 transactions consuming 113,265 gas, while the losing searcher issued 43 transactions consuming 33,547 gas. Further details on which MEV transactions are value-creating and which are value-diverting are explained in Section 5.
@@ 3 Related Survey
Several studies have examined different types of MEV transactions and their mitigation strategies within the DeFi ecosystem [7, 8, 6, 5, 9]. Li et al. [7] analyzed vulnerabilities, attacks, and security optimizations in DeFi across different blockchain layers, noting that the consensus layer is susceptible to front-running and sandwich transactions that contribute to MEV. However, the survey lacks a thorough analysis of MEV transactions and does not address detection approaches or mitigation strategies.
Alipanahloo et al. [8] categorized MEV mitigation strategies based on their potential to reduce or democratize MEV. They highlighted that fair transaction ordering policies, encrypted mempool and private transactions, and smart contract-based strategies are effective in reducing MEV, while the Proposer-Builder Separation (PBS) approach serves to democratize MEV. The study also assessed the impact of these strategies on network performance and implementation challenges. The authors find that mempool privacy through delay encryption and PBS strategies show promise in countering unfair MEV opportunities. Mohan and Khezr [6] presented a temporal evolution of Ethereum protocols driven by the underlying need for MEV mitigation. However, their focus is limited to mitigation strategies based on private transactions. However, these studies [6, 8] do not consider MEV transactions and detection approaches.
Poux et al. [5] provided a categorization of MEV transactions based on fairness and analyzed their impact on the network. They highlighted that liquidation and arbitrage transactions have a positive impact, whereas front-running, sandwich, and time-bandit transactions negatively affect the system. Furthermore, their study briefly discussed the private transactions-based MEV mitigation strategies. Rasheed et al. [9] analyzed front-running, back-running, and sandwich MEV transactions along with the evolution of private transactions-based MEV mitigation strategies. While these studies [5, 9] cover various MEV transaction types and private transactions-based mitigation strategies, they do not analyze different types of MEV within each transaction type, nor do they explore other categories of mitigation strategies.
Table I summarizes the focus of each study. As shown, most surveys concentrate on MEV mitigation strategies [6, 8, 5, 9], with [7, 5, 9] covering a few categories of MEV transactions. To our knowledge, no survey comprehensively addresses MEV from the perspectives of transaction types, detection approaches, mitigation strategies, and simulation methods. This survey aims to fill this gap.
@@ 4.1 Literature Retrieval
This step involves selecting databases, formulating search strings, and retrieving articles. To conduct a comprehensive survey, we queried multiple databases, including the ACM Digital Library, IEEE Xplore, ScienceDirect, Scopus, Springer, Taylor & Francis, and Web of Science. We developed search strings focusing on ”Maximal/Miner Extractable Value” and ”Ethereum/Smart Contract” to extract relevant articles. Table II presents the search strings used for each database. The search was limited to publications up to and including May 20, 2024. Following the initial database search, we performed an additional manual search on Google Scholar to identify articles that were not extracted via database search. This approach ensures an exhaustive exploration of the literature.
| Database | Search string |
| ACM | Abstract:((”MEV” OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Ethereum” OR ”Smart Contract*” OR ”Smart-Contract*”)) OR Title:((”MEV” OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Ethereum” OR ”Smart Contract*” OR ”Smart-Contract*”)) OR Keyword:((”MEV” OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Ethereum” OR ”Smart Contract*” OR ”Smart-Contract*”)) |
| IEEE Xplore |
(”Document Title”: MEV OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Document Title”: Ethereum OR ”Smart Contract* OR ”Smart-Contract*”)
(”Abstract”: MEV OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Abstract”: Ethereum OR ”Smart Contract* OR ”Smart-Contract*”) (”Author Keywords”: MEV OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Author Keywords”: Ethereum OR ”Smart Contract* OR ”Smart-Contract*”) |
| Science Direct |
(”MEV” OR ”Maximal extractable value” OR ”Maximal-extractable-value” OR ”Miner extractable value” OR ”Miner-extractable-value”) AND (Ethereum OR ”Smart contract” OR ”smart-contract”)
(”MEV” OR ”Maximal extractable values” OR ”Maximal-extractable-values” OR ”Miner extractable values” OR ”Miner-extractable-values”) AND (Ethereum OR ”Smart contracts” OR ”smart-contracts”) |
| Scopus | TITLE-ABS-KEY ((”MEV” OR ”Maximal-Extractable-value*” OR ”Maximal Extractable value*” OR ”Miner-Extractable-value*” OR ”Miner Extractable value*”) AND (”Ethereum” OR ”Smart Contract*” OR ”Smart-Contract*”)) |
| Springer | ”MEV” OR ”Maximal extractable value*” OR ”Maximal-extractable-value*” OR ”Miner extractable value*” OR ”Miner-extractable-value*” AND (”Ethereum” OR ”Smart contract*” OR ”smart-contract*”) |
| Taylor & Francis |
[[Publication Title: ”mev”] OR [Publication Title: ”maximal extractable value*”] OR [Publication Title: ”maximal-extractable-value*”] OR [Publication Title: ”miner extractable value*”] OR [Publication Title: ”miner-extractable-value*”]] AND [[Publication Title: ethereum] OR [Publication Title: ”smart contract*”] OR [Publication Title: ”smart-contract*”]]
[[Keywords: ”mev”] OR [Keywords: ”maximal extractable value*”] OR [Keywords: ”maximal-extractable-value*”] OR [Keywords: ”miner extractable value*”] OR [Keywords: ”miner-extractable-value*”]] AND [[Keywords: ethereum] OR [Keywords: ”smart contract*”] OR [Keywords: ”smart-contract*”]] [[Abstract: ”mev”] OR [Abstract: ”maximal extractable value*”] OR [Abstract: ”maximal-extractable-value*”] OR [Abstract: ”miner extractable value*”] OR [Abstract: ”miner-extractable-value*”]] AND [[Abstract: ethereum] OR [Abstract: ”smart contract*”] OR [Abstract: ”smart-contract*”]] |
| Web of Science | (AB=((”MEV” OR ”Maximal extractable value*” OR ”Maximal-extractable-value*” OR ”Miner extractable value*” OR ”Miner-extractable-value*”) AND (Ethereum OR ”Smart contract*” OR ”smart-contract*”))) OR (TI=((”MEV” OR ”Maximal extractable value*” OR ”Maximal-extractable-value*” OR ”Miner extractable value*” OR ”Miner-extractable-value*”) AND (Ethereum OR ”Smart contract*” OR ”smart-contract*”))) OR (AK=((”MEV” OR ”Maximal extractable value*” OR ”Maximal-extractable-value*” OR ”Miner extractable value*” OR ”Miner-extractable-value*”) AND (Ethereum OR ”Smart contract*” OR ”smart-contract*”))) |

### v2

@@ Abstract
Decentralized Exchanges (DEXes) are one of the prominent applications of DeFi, enabling users to directly swap tokens using smart contracts [2]. Most DEXes operate on the Ethereum blockchain to facilitate trustless and automated transactions. However, in the Ethereum network, transactions are stored in a public mempool before they are included in a block. This creates profitable opportunities for network participants, referred to as searchers, who can submit new transactions by observing pending financial transactions to gain additional revenue. These searchers maximize their profits by manipulating gas prices to influence the order of their transactions within the block. This additional value extracted from the blockchain network is termed Maximal Extractable Value (MEV) [3]. MEV searchers can be the block producers (i.e., miners, validators, or block proposers), other network participants, or bots. However, block producers have a unique advantage, as they can include transactions in blocks without paying high gas prices [4]. Furthermore, the transition from Proof-of-Work (PoW) to Proof-of-Stake (PoS) in Ethereum has significantly reduced block producer rewards, further luring block producers to engage in MEV activities [5].
@@
Several surveys have explored various aspects of MEV, including transaction types and mitigation strategies [8, 9, 7, 6, 10, 11]. In particular, [8] classifies MEV transactions, while [9, 7] emphasize on mitigation techniques. In contrast, [6, 10, 11] cover both transaction types and mitigation strategies. However, none have thoroughly examined all transaction types, detection approaches, and mitigation strategies. This comprehensive survey addresses this gap by presenting a novel taxonomy of MEV transaction types supported by real-world examples from the Ethereum network. It also explores MEV detection approaches and mitigation strategies and examines various simulation and extraction methods. The main contributions of this survey are as follows.
We introduce a novel and comprehensive taxonomy of MEV transactions, supported by real Ethereum transaction examples for each identified MEV type. The taxonomy distinguishes between value-diverting MEV transactions, which can lead to financial loss and network instability, and value-creating MEV transactions, which can stabilize markets or enhance efficiency. This distinction provides clarity on the dual nature of MEV activities and lays the foundation for discussions on detection and mitigation strategies.
@@
We explore existing MEV simulation frameworks and extraction methods, offering a comprehensive comparison of how these methods model and replicate real-world MEV scenarios.
We critically examine the major challenges in MEV detection, mitigation, and simulation, such as centralization, latency, cross-domain MEV, multi-address MEV, and simulation environment. Furthermore, we discuss potential solutions paving the way for a more secure and efficient DeFi system.
@@ 2.1.1 EIP-1599
2.2 Decentralized Finance (DeFi) Applications
DeFi applications offer a wide range of financial services through decentralized protocols [16, 11]. Below, we provide an overview of the widely used DeFi applications. This will help in understanding how various MEV transactions can impact these applications.
DEXes are blockchain-based trading platforms that eliminate the need for a centralized operator. Unlike traditional centralized exchanges (CEXes), where an exchange operator maintains an order book and facilitates trade execution [17], DEXes use smart contracts to enable direct peer-to-peer asset trading [2]. The trading mechanisms in DEXes can be either order book-based, where orders are matched manually (as in EtherDelta) or automatically by a smart contract (as in IDEX) [13], or based on Automated Market Makers (AMMs), which use smart contracts to create liquidity pools. In AMMs, users trade assets directly against the pool, with prices determined algorithmically based on the new ratio of assets in the pool. Xu et al. [2] provides a detailed analysis of AMM pricing models. Larger asset trades result in greater price shifts, leading to slippage, where traders receive a less favorable price than expected due to the impact of their trade on the pool’s balance.
Lending and borrowing platforms in DeFi allow users to lend and borrow assets from lending pools. Lenders earn interest paid by borrowers, who must provide collateral as security. The value of the collateral must exceed a minimum ratio, which varies depending on the lending protocol and asset (e.g., 150% for certain assets). If the collateral’s value falls below the liquidation ratio due to market volatility, liquidators can repay part of the loan in exchange for acquiring the collateral at a discounted price. This process, known as liquidation, ensures that lenders are repaid and prevents bad debt accumulation. Borrowers can avoid liquidation by depositing additional collateral or repaying part of the loan to maintain a safe ratio.
Yield farming enables users to generate returns by depositing assets into liquidity pools or lending protocols in exchange for interest and trading fees. On the other hand, Margin trading allows users to borrow funds to leverage their trades, with smart contracts automating the borrowing, lending, and liquidation processes. Derivatives in DeFi enable users to trade financial contracts (derivatives) that are linked to underlying crypto assets. Asset management platforms allow users to optimize and manage their cryptocurrency holdings using tools for portfolio management, risk management, and yield optimization. In prediction markets, users can place bets on the outcomes of future events, such as elections, sports matches, or financial market movements, with decentralized oracles ensuring trustless and transparent verification.
Tokenization in DeFi has made it possible to convert real-world physical assets like real estate and commodities into digital tokens, enabling their trade within the DeFi ecosystems. Domain name protocols allows the creation of censorship-resistant decentralized domains to host websites or services that are not subject to traditional censorship or regulation. To enhance the security of DeFi protocols, some platforms reward users for identifying vulnerabilities in smart contracts. Lastly, Gambling and lottery platforms within DeFi offer transparent and trustless participation gambles and lotteries, with outcomes governed by smart contracts.
@@ 2.3 Maximal Extractable Value (MEV)
Publicly visible transactions on Ethereum enable searchers to extract MEV through the insertion, removal, or reordering of transactions [13]. According to Ethereum’s documentation444https://ethereum.org/en/developers/docs/mev/, accessed on 10 February 2024: Maximal extractable value (MEV) refers to the maximum value that can be extracted from block production in excess of the standard block reward and gas fees by including, excluding, and changing the order of transactions in a block.
@@
Before the Merge555The update that transitioned Ethereum’s consensus protocol from Proof-of-Work (PoW) to Proof-of-Stake (PoS) to improve scalability, security, and energy efficiency, Ethereum was based on the PoW consensus protocol, where block producers were known as miners. The value extracted from block production was thus initially referred to as Miner Extractable Value. After the Merge in September 2022, Ethereum transitioned to the PoS consensus protocol, with validators taking the place of miners. Additionally, a searcher does not necessarily have to be the block producer; they can also be non-miner users or bots. Consequently, the term ”Miner” in MEV was replaced with ”Maximal” to reflect this change. MEV is sometimes also referred to as ”Blockchain Extractable Value” [18]. Furthermore, the Flashbots project refers to the actual extracted values due to MEV opportunities as ”Realized Extractable Value”666https://writings.flashbots.net/quantifying-rev, accessed on 05 September 2024. Judmayer et al. [19] describe the potential or future values that could be extracted if a sequence of transactions eventually becomes a part of the blockchain (with some probability) as ”Expected Extracted Value”.
2.3.1 MEV Transactions
The main aspects of MEV include the actions performed to extract MEV (i.e., insertion, removal, or reordering), the network participants who act as searchers, and the nature of MEV transactions (i.e., single or multi-block). Based on these aspects, we define MEV transactions as transactions inserted by searchers (block producers, bots, or other network participants) to gain a financial advantage in DeFi by exploiting victim transactions777A transaction in the public mempool that presents an opportunity for a searcher to generate profits. or market fluctuations. These transactions can span multiple blocks, with their order determined through gas price adjustments and/or private tips. Searchers can temporarily censor a victim transaction by executing multiple MEV transactions. It is important to note that not all MEV transactions are inherently harmful. Some enhance market efficiency, while others lead to financial losses, network congestion, or blockchain instability.
@@ 2.3.1 MEV Transactions
For instance, a searcher may identify a price difference for the same asset across different DeFi protocols due to trading activities. The searcher performs an MEV transaction to profit from this price difference, which helps stabilize asset prices across these platforms. Such transactions are known as ”value-creating” MEV transactions, as they generate new value [7]. These transactions are generally regarded as beneficial rather than harmful.
@@
On the other hand, some searchers monitor the mempool for MEV-related transactions and attempt to preempt them, or might deliberately execute a transaction that generates an MEV opportunity and later execute another transaction to capture that MEV. For example, a searcher might detect a potential victim transaction aimed at exploiting an arbitrage opportunity, copy it, and then resubmit it with a higher gas fee to capture the MEV before the original transaction is executed. Such transactions are referred to as ”value-diverting” MEV transactions because they divert value from victims to the searcher [7]. Such transactions can be considered harmful, as they negatively impact other users. Value-diverting transactions have a detrimental impact on the DeFi system, especially when multiple searchers engage in open-bid first-price888In a first-price auction, the winning searcher pays their bid amount. on-chain auctions to capture MEV. These competitive auctions, known as Priority Gas Auctions (PGAs) [13], force searchers to pay significantly higher gas fees to prioritize their transactions. This bidding war increases the average gas price across the network, discouraging honest users from utilizing the platform. Additionally, searchers often submit multiple redundant transactions with progressively higher gas fees to maintain their position in the auction, leading to network congestion and increased latency [7]. Daian et al. [13] illustrates an example of a PGA between two searchers, where the winning searcher issued 42 transactions consuming 113,265 gas, while the losing searcher issued 43 transactions consuming 33,547 gas. Further details on which MEV transactions are value-creating and which are value-diverting are explained in Section 5.
@@ 3 Related Survey
Several studies have examined different types of MEV transactions and their mitigation strategies within the DeFi ecosystem [8, 9, 7, 6, 10, 11]. Li et al. [8] analyzed vulnerabilities, attacks, and security optimizations in DeFi across different blockchain layers, noting that the consensus layer is susceptible to front-running and sandwich transactions that contribute to MEV. However, the survey lacks a thorough analysis of MEV transactions and does not address detection approaches or mitigation strategies.
Alipanahloo et al. [9] categorized MEV mitigation strategies based on their potential to reduce or democratize MEV. They highlighted that fair transaction ordering policies, encrypted mempool and private transactions, and smart contract-based strategies are effective in reducing MEV, while the Proposer-Builder Separation (PBS) approach serves to democratize MEV. The study also assessed the impact of these strategies on network performance and implementation challenges. The authors find that mempool privacy through delay encryption and PBS strategies show promise in countering unfair MEV opportunities. Mohan and Khezr [7] presented a temporal evolution of Ethereum protocols driven by the underlying need for MEV mitigation. However, their focus is limited to mitigation strategies based on private transactions. However, these studies [7, 9] do not consider MEV transactions, detection approaches, and simulation methods.
Poux et al. [6] provided a categorization of MEV transactions based on fairness and analyzed their impact on the network. They highlighted that liquidation and arbitrage transactions have a positive impact, whereas front-running, sandwich, and time-bandit transactions negatively affect the system. Furthermore, their study briefly discussed the private transactions-based MEV mitigation strategies. Rasheed et al. [10] analyzed front-running, back-running, and sandwich MEV transactions along with the evolution of private transactions-based MEV mitigation strategies. Lastly, Gramlich et al. [11] provided details on front-running, back-running, sandwich, arbitrage, and liquidation transactions, along with various countermeasures for mitigating MEV. While these studies [6, 10, 11] do not analyze the different types of MEV within each transaction type, nor do they explore detection approaches and simulation methods.
Table I summarizes the focus of each study. As shown, most surveys concentrate on MEV mitigation strategies [7, 9, 6, 10, 11], with [8, 6, 10, 11] covering a few categories of MEV transactions. To our knowledge, no survey comprehensively addresses MEV from the perspectives of transaction types, detection approaches, mitigation strategies, and simulation methods. This survey aims to fill this gap.
@@ 4.1 Literature Retrieval
This step involves selecting databases, formulating search strings, and retrieving articles. To conduct a comprehensive survey, we queried multiple databases, including the ACM Digital Library, IEEE Xplore, ScienceDirect, Scopus, Springer, Taylor & Francis, and Web of Science. We developed search strings focusing on ”Maximal/Miner Extractable Value” and ”Ethereum/Smart Contract” to extract relevant articles. The initial search was limited to publications up to and including May 20, 2024. Following the initial database search, we performed an additional manual search on Google Scholar to identify articles that were not extracted via database search. This approach ensures an exhaustive exploration of the literature.